[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sun Jul 16, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Mon Jul 17 02:31:26 UTC 2017


HSU Shift Log for Sun Jul 16 2017 22:31:20 GMT-0400 (EDT)

Shift: R

Forecaster: E Blake

Technical Issues Summary:

ATCF issues:

N-AWIPS issues: The GOES-16 RGB imagery that Monica set up late last week  
is not updating. That might have been turned off as a fix for the disk  
filling up, but wanted to mention it.

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170717/5c962332/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list